# Positive microlocal holonomies are globally regular

### v1

@@ 2.5.4. Legendrian links $\Lambda_{\beta}$
Example 2.5.2 generalizes to Legendrian links of the form $\Lambda=\Lambda_{\beta}$, $\beta$ a positive braid in $n$-strands. By using Equation (2.4), we may assume that the set $\mathfrak{t}$ of basepoints has exactly one basepoint per component. Set $G:=\operatorname{GL}_{1}^{|\mathfrak{t}|}$, $H:=\operatorname{GL}_{1}$ and identify $H$ with the diagonal subgroup $\Delta(H)\subset G$. By [CW24, Section 4.1], see also [CGG+22, Corollary 3.7], there exists a smooth affine variety $X_{\beta}$ such that
@@
where $H$ acts trivially on $X_{\beta}$. That said, as illustrated in Example 2.5.2, the $(G/H)$-action on $X_{\beta}$ is not free and the quotient $\mathfrak{M}_{1}({\Lambda_{\beta}})$ typically has non-trivial stabilizers (and different from $H$) at some points. As in Example 2.5.1, $X_{\beta}$ is closely related to a braid variety: specifically, per Equation (2.4), the braid variety $X(\beta)$ is isomorphic to $X_{\beta\delta(\beta)}\times\operatorname{GL}_{1}^{n-|\mathfrak{t}|}$, cf. [CL23, Proposition 5.8]. An explicit description of the basepoint action follows from [CGGS20, Section 2.2], cf. also [CN22, Section 5]. $\Box$
@@ 3.1. The Kashiwara-Schapira stack
[…]
@@
First, we define an $\infty$-sheaf of dg-categories in the cotangent bundle $T^{*}M$ of a smooth manifold $M$. By definition, a subset $\widehat{\Lambda}\subset T^{*}M$ is conic if it is invariant under the following scaling $\mathbb{R}_{+}$-action: +×$\mathbb{R}_{+}\times T^{*}M\to T^{*}M,(s;x,\xi)\to(x,s\xi)$. A typical example of $\widehat{\Lambda}$ will be the union of the Lagrangian cone over a Legendrian submanifold $\Lambda\subset T^{*}_{\infty}M$ and the zero section $M\subset T^{*}M$. In particular, $\widehat{\Lambda}$ might be singular.
@@
Note that for any (conic) open set $\widehat{\Omega}\subset T^{*}M$ with $\widehat{\Omega}\cap\widehat{\Lambda}=\emptyset$ we have a dg-equivalence $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}^{\text{pre}}(\widehat{\Omega})\cong 0$, where the right hand side is understood as the one-object category with the zero vector space as its endomorphisms. Therefore $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}^{\text{pre}}$ is supported on $\widehat{\Lambda}$. Hence, $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}$ is also supported on $\widehat{\Lambda}$. We often abuse notation and still denote the restriction of this sheaf to $\widehat{\Lambda}$ by $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}:\operatorname{Op}(\widehat{\Lambda})^{op}\longrightarrow\mbox{dg-Cat}_{k}$.
@@
$\hskip 31.0pt\operatorname{Sh}_{\Lambda}(M)\hookrightarrow\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}M\backslash\Omega)}(M)\to\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}M\backslash\Omega)}(M)/\operatorname{Sh}_{T^{*}_{\infty}M\backslash\Omega}(M)=\operatorname{\operatorname{\mu Sh}}_{\Lambda}^{\text{pre}}(\Omega).\hskip 21.0pt\Box$
@@ 3.1.1. Comments and examples related to the Kashiwara-Schapira stack
The literature contains alternative descriptions for such sheaves $\operatorname{\operatorname{\mu Sh}}_{\Lambda}$, see e.g. [Gui23, Section 10] or [Nad16, Section 3.4]. For comparison to the former, consider $\Lambda\subset T^{*}_{\infty}M$ and an open subset $\Lambda_{0}\subset\Lambda$. Using the notation $\operatorname{Sh}_{(\Lambda_{0})}(M)=\bigcup_{\Lambda_{0}\subset\Omega}\operatorname{Sh}_{\Lambda_{0}\cup T^{*}_{\infty}M\backslash\Omega}(M)$ from [Gui23, Section 10], we have
@@
Since the restriction $r^{*}:\operatorname{Sh}(M)\to\operatorname{Sh}(B)$ preserves limits and colimits (whose left adjoint is $r_{!}$ and right adjoint is $r_{*}$), the restriction on the full subcategories $\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}M\backslash\Omega)}(M)$ $\to\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}B\backslash\Omega)}(B)$ also preserves limits and colimits and admits left and right adjoints. (Note that [Kuo23, Thm. 1.2] provides explicit descriptions of these functors.) Using the (left) adjoint functor $(r^{*})^{\ell}$, one can show that, given $\mathscr{F}\in\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}M\backslash\Omega)}(M)$, the map $\mathscr{F}\to r^{*\ell}r^{*}\mathscr{F}$ is an isomorphism in $\Omega$, and, similarly, given $\mathscr{F}_{B}\in\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}B\backslash\Omega)}(B)$, $r^{*}r^{*\ell}\mathscr{F}_{B}\to\mathscr{F}_{B}$ is also an isomorphism in $\Omega$, i.e. the mapping cones have singular supports in $T^{*}_{\infty}M\setminus\Omega$ and $T^{*}_{\infty}B\setminus\Omega$. Thus we obtain the equivalence between the definition in [Nad16, Section 3.4] and Definition 3.3 above. $\Box$
@@
(Continued) If the projection $\pi:\Lambda\to M$ is finite-to-one, [Nad16, Section 3.4] and [JT24, Section 3.9–11] use the definition
$\operatorname{\operatorname{\mu Sh}}_{\Lambda}^{\text{pre}}(\Omega)=\operatorname{Sh}_{\Lambda}(B)/\mathrm{Loc}(B).$
The equivalence between this and the above ones is an application of the refined microlocal cut-off lemma, cf.  [KS90, Prop. 6.1.4] and [Gui23, Lem. 10.2.5]. Namely, using the microlocal cut-off functor $\iota^{\ell}$ it follows that, given $\mathscr{F}_{B}\in\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}B\backslash\Omega)}(B)$, the map $\mathscr{F}_{B}\to\iota^{\ell}\mathscr{F}_{B}$ is an isomorphism on $\Omega$, and similarly, given $\mathscr{F}_{B}^{\prime}\in\operatorname{Sh}_{\Lambda}(B)$, the map $\iota^{\ell}\mathscr{F}_{B}^{\prime}\to\mathscr{F}_{B}^{\prime}$ is an isomorphism. (See also the arguments in [KL22, Cor. 4.32].) This implies the equivalence of this definition and the above one.
As in [Nad16, Section 3.4], one can even define the sheaf of categories locally by
@@
The isomorphism between $\operatorname{\operatorname{\mu Sh}}_{\Lambda}$ and $\operatorname{\operatorname{\mu Sh}}^{\text{pre}}_{\Lambda}$ on such small open balls follows from the constructibility of the internal Hom in $\operatorname{Sh}_{\Lambda}(B)$ when $\Lambda\subset T^{*}_{\infty}M$ is subanalytic Legendrian [KS90, Prop. 8.4.6], which implies local constancy of the Hom when we restrict to smaller and smaller neighbourhoods around a point [KS90, Lem. 8.4.7]. See also [KL22, Lem. 4.22 & Cor. 4.32].
$\Box$
@@
This Lagrangian skeleton $\mathbb{K}_{n}$ corresponds to the case that $\Lambda\subset T^{*}_{\infty}M$ is a set of $n$ points, and $\widehat{\Lambda}=\mathbb{K}_{n}(M)$ is given by the union of the (Lagrangian) cone over $\Lambda$ union the zero section $M\subset T^{*}M$. We use the cases $M=S^{1}$ and $M=[0,1]$ in Subsection 4.2, cf. Lemmas 4.2 and 4.5.
@@
Let $\Lambda\subset T^{*}_{\infty}\mathbb{R}^{2}$ be a Legendrian link and choose $\widehat{\Lambda}\subset T^{*}\mathbb{R}^{2}$ to be the union of the Lagrangian cone over $\Lambda$ and all the bounded regions of $\mathbb{R}^{2}\setminus\pi(\Lambda)$, where $\pi:T^{*}M\longrightarrow M$ is the natural projection. For any $U\subset\mathbb{R}^{2}$, we have $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}^{\text{pre}}(T^{*}U)=\operatorname{Sh}_{\widehat{\Lambda}}(U)$. Now, $\operatorname{Sh}_{\widehat{\Lambda}}$ forms a sheaf of categories and the universal property of sheafification implies that $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}(T^{*}U)=\operatorname{Sh}_{\widehat{\Lambda}}(U)$. ((((See also [JT24, Appendix B].)))) In particular, since $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}$ is supported on $\widehat{\Lambda}$, global sections of the Kashiwara–Schapira stack are
@@
where the subscript 00 denotes the full subcategory of sheaves with bounded support in $\mathbb{R}^{2}$. $\Box$

### v2

@@ 2.5.4. Legendrian links $\Lambda_{\beta}$
Example 2.5.2 generalizes to Legendrian links of the form $\Lambda=\Lambda_{\beta}$, $\beta$ a positive braid in $n$-strands. By using Equation (2.5), we may assume that the set $\mathfrak{t}$ of basepoints has exactly one basepoint per component. Set $G:=\operatorname{GL}_{1}^{|\mathfrak{t}|}$, $H:=\operatorname{GL}_{1}$ and identify $H$ with the diagonal subgroup $\Delta(H)\subset G$. By [CW24, Section 4.1], see also [CGG+25, Corollary 3.7], there exists a smooth affine variety $X_{\beta}$ such that
@@
where $H$ acts trivially on $X_{\beta}$. That said, as illustrated in Example 2.5.2, the $(G/H)$-action on $X_{\beta}$ is not free and the quotient $\mathfrak{M}_{1}({\Lambda_{\beta}})$ typically has non-trivial stabilizers (and different from $H$) at some points. As in Example 2.5.1, $X_{\beta}$ is closely related to a braid variety: specifically, per Equation (2.5), the braid variety $X(\beta)$ is isomorphic to $X_{\beta\delta(\beta)}\times\operatorname{GL}_{1}^{n-|\mathfrak{t}|}$, cf. [CW24, Section 4.1] or [CL23, Proposition 5.8]. An explicit description of the basepoint action follows from [CGG+24, Section 2.2], cf. also [CN22, Section 5] from the perspective of Legendrian contact dg algebra. $\Box$
@@ 3.1. The Kashiwara-Schapira stack
The category $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})$ is the global sections of a sheaf of dg-categories, known as the Kashiwara-Schapira stack. This subsection reviews such stack and its properties. Let $X$ be a Weinstein sector with subanalytic Lagrangian skeleton $\mathbb{L}$. In this work, the expression sheaves of dg-categories precisely means $\infty$-sheaves, cf. [LUR09, Def. 7.3.3.1], valued in the $\infty$-category $\mbox{dg-Cat}_{k}$ of well generated dg-categories, cf. [TAB09] (we will see later in the proof of Theorem 3.16 that these sheaves on Lagrangian skeleta are in fact valued in the $\infty$-category of compactly generated dg-categories).
@@
First, we define an $\infty$-sheaf of dg-categories in the cotangent bundle $T^{*}M$ of a smooth manifold $M$. By definition, a subset $\widehat{\Lambda}\subset T^{*}M$ is conic if it is invariant under the following scaling $\mathbb{R}_{+}$-action: $\mathbb{R}_{+}\times T^{*}M\to T^{*}M,(s;x,\xi)\to(x,s\xi)$. A typical example of $\widehat{\Lambda}$ will be the union of the Lagrangian cone over a Legendrian submanifold $\Lambda\subset T^{*}_{\infty}M$ and the zero section $M\subset T^{*}M$. In particular, $\widehat{\Lambda}$ might be singular.
@@
Note that for any (conic) open set $\widehat{\Omega}\subset T^{*}M$ with $\widehat{\Omega}\cap\widehat{\Lambda}$$\emptyset$ we have a dg-equivalence $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}^{\text{pre}}(\widehat{\Omega})\cong 0$, where the right hand side is understood as the one-object category with the zero vector space as its endomorphisms. Therefore $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}^{\text{pre}}$ is supported on $\widehat{\Lambda}$. Hence, $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}$ is also supported on $\widehat{\Lambda}$. We often abuse notation and still denote the restriction of this sheaf to $\widehat{\Lambda}$ by $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}:\operatorname{Op}(\widehat{\Lambda})^{op}\longrightarrow\mbox{dg-Cat}_{k}$.
@@
$\hskip 42.0pt\operatorname{Sh}_{\Lambda}(M)\hookrightarrow\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}M\backslash\Omega)}(M)\to\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}M\backslash\Omega)}(M)/\operatorname{Sh}_{T^{*}_{\infty}M\backslash\Omega}(M)=\operatorname{\operatorname{\mu Sh}}_{\Lambda}^{\text{pre}}(\Omega).\hskip 37.0pt\Box$
@@ 3.1.1. Comments and examples related to the Kashiwara-Schapira stack
The literature contains alternative descriptions for such sheaves $\operatorname{\operatorname{\mu Sh}}_{\Lambda}$, see e.g. [GUI23, Section 10] or [NAD16, Section 3.4]. For comparison to the former (defined for smooth Legendrians), consider $\Lambda\subset T^{*}_{\infty}M$ and an open subset $\Lambda_{0}\subset\Lambda$. Using the notation $\operatorname{Sh}_{(\Lambda_{0})}(M)=\bigcup_{\Lambda_{0}\subset\Omega}\operatorname{Sh}_{\Lambda_{0}\cup T^{*}_{\infty}M\backslash\Omega}(M)$ from [GUI23, Section 10], we can show by [GUI23, Lemma 10.2.2] that
@@
Since the restriction $r^{*}:\operatorname{Sh}(M)\to\operatorname{Sh}(B)$ admits left adjoint $r_{!}$ and right adjoint $r_{*}$, and the inclusion $\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}M\backslash\Omega)}(M)\hookrightarrow\operatorname{Sh}(M)$ admits left and a right adjoints by [KUO23, Thm. 1.2], we know that the restriction on the full subcategories $r^{*}:\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}M\backslash\Omega)}(M)\to\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}B\backslash\Omega)}(B)$ also admits left and right adjoints. Using the formula of the left adjoint functor $r^{*\ell}$ in [KUO23, Thm. 1.2], one can show that, given $\mathscr{F}\in\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}M\backslash\Omega)}(M)$, the map $\mathscr{F}\to r^{*\ell}r^{*}\mathscr{F}$ is an isomorphism in $\Omega$ (i.e. the mapping cone has singular support in $T^{*}_{\infty}M\setminus\Omega$), and, similarly, given $\mathscr{F}_{B}\in\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}B\backslash\Omega)}(B)$, the map $r^{*}r^{*\ell}\mathscr{F}_{B}\to\mathscr{F}_{B}$ is also an isomorphism in $\Omega$. Thus we obtain the equivalence between the definition in [NAD16, Section 3.4] and Definition 3.3 above. $\Box$
If the projection $\pi:\Lambda\to M$ is finite-to-one, we can connect our definition with [NAD16, Section 3.4] and [JT24, Section 3.9–11]:
Lemma 3.6.
For $\Lambda\subset T^{*}_{\infty}M$, if the projection $\pi:\Lambda\to M$ is finite-to-one, then for any sufficiently small open ball $B\subset M$ and $\Omega\subset T^{*}_{\infty}B$,
$\operatorname{\operatorname{\mu Sh}}_{\Lambda}^{\text{pre}}(\Omega)\cong\operatorname{Sh}_{\Lambda}(B)/\mathrm{Loc}(B).$
Proof.
Per Remark 3.5, it suffices to show that $\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}B\setminus\Omega)}(B)/\operatorname{Sh}_{T^{*}_{\infty}B\setminus\Omega}(B)\cong\operatorname{Sh}_{\Lambda}(B)/\mathrm{Loc}(B)$. This is an application of the refined microlocal cut-off lemma, cf.  [KS90, Prop. 6.1.4] and [GUI23, Lem. 10.2.5]: The inclusion $\iota_{*}:\operatorname{Sh}_{\Lambda}(B)\hookrightarrow\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}B\backslash\Omega)}(B)$ admits a left adjoint $\iota^{*}$ called the (refined) microlocal cut-off functor. Given $\mathscr{F}_{B}^{\prime}\in\operatorname{Sh}_{\Lambda}(B)$, the map $\iota^{*}\iota_{*}\mathscr{F}_{B}^{\prime}\to\mathscr{F}_{B}^{\prime}$ is always an isomorphism, and given $\mathscr{F}_{B}\in\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}B\backslash\Omega)}(B)$, by the refined microlocal cut-off lemma, $\mathscr{F}_{B}\to\iota_{*}\iota^{*}\mathscr{F}_{B}$ is also an isomorphism on $\Omega$. This completes the proof. (See also [KL22, Lem. 4.28 & Cor. 4.32].) ∎
Lemma 3.7.
For $\Lambda\subset T^{*}_{\infty}M$, if the projection $\pi:\Lambda\to M$ is finite-to-one, then for any sufficiently small open ball $B\subset M$, $\Omega\subset T^{*}_{\infty}B$,
$\operatorname{\operatorname{\mu Sh}}_{\Lambda}^{\text{pre}}(\Omega)\cong\operatorname{\operatorname{\mu Sh}}_{\Lambda}^{\text{pre}}(\Omega)\cong\operatorname{Sh}_{\Lambda}(B)/\mathrm{Loc}(B).$
Proof.
We show that for sufficiently small open balls $B^{\prime}\subset B$, the restriction functor
$r^{*}:\operatorname{Sh}_{\Lambda}(B)/\mathrm{Loc}(B)\longrightarrow\operatorname{Sh}_{\Lambda}(B^{\prime})/\mathrm{Loc}(B^{\prime})$
induces an equivalence. Then, by Lemma 3.6, it follows that for $\Omega^{\prime}\subset\Omega$, the restriction functor $\operatorname{\operatorname{\mu Sh}}_{\Lambda}^{\text{pre}}(\Omega)\longrightarrow\operatorname{\operatorname{\mu Sh}}_{\Lambda}^{\text{pre}}(\Omega^{\prime})$ also induces an equivalence. Hence for the $\infty$-sheafification we also have
$\operatorname{\operatorname{\mu Sh}}_{\Lambda}^{\text{pre}}(\Omega)\cong\operatorname{\operatorname{\mu Sh}}_{\Lambda}^{\text{pre}}(\Omega)\cong\operatorname{Sh}_{\Lambda}(B)/\mathrm{Loc}(B).$
For the restriction $r^{*}:\operatorname{Sh}_{\Lambda}(B)/\mathrm{Loc}(B)\longrightarrow\operatorname{Sh}_{\Lambda}(B^{\prime})/\mathrm{Loc}(B^{\prime})$, it is fully faithful by the constructibility of the internal Hom in $\operatorname{Sh}_{\Lambda}(B)$ when $\Lambda\subset T^{*}_{\infty}M$ is subanalytic Legendrian [KS90, Prop. 8.4.6], which implies local constancy of the Hom when we restrict to smaller and smaller neighbourhoods around a point [KS90, Lem. 8.4.7]. Since the front projection of $\Lambda$ defines a Whitney stratification, such that the inclusions of the strata in $B^{\prime}$ to $B$ are homotopy equivalences [MAT12, Sec. 7 & 8]. Hence any constructible sheaf in $\operatorname{Sh}_{\Lambda}(B^{\prime})/\mathrm{Loc}(B^{\prime})$ extends to a constructible sheaf in $B$ which is contained in $\operatorname{Sh}_{\Lambda}(B)/\mathrm{Loc}(B)$ by checking the condition [GPS24, Cor. 4.23]. This shows the restriction $r^{*}$ is essential surjective. (See also [KL22, Lem. 4.28 & Cor. 4.32].) ∎
@@
In fact, in [NAD16, Section 3.4], the sheaf of categories is directly defined as a constructible sheaf of categories such that
$\operatorname{\operatorname{\mu Sh}}_{\Lambda}(\Omega)=\operatorname{Sh}_{\Lambda\cup(T^{*}_{\infty}B\backslash\Omega)}(B)/\operatorname{Sh}_{T^{*}_{\infty}B\backslash\Omega}(B).$
and it is mentioned that when the projection $\pi:\Lambda\to M$ is finite-to-one, one can furthermore write
@@
The above lemma shows that our definition agrees with the description there. ∎
@@
This Lagrangian skeleton $\mathbb{K}_{n}$ corresponds to the case that $\Lambda\subset T^{*}_{\infty}M$ is a set of $n$ points, and $\widehat{\Lambda}=\mathbb{K}_{n}(M)$ is given by the union of the (Lagrangian) cone over $\Lambda$ union the zero section $M\subset T^{*}M$. We use the cases $M=S^{1}$ and $M=[0,1]$ in Subsection 4.2, cf. Lemmas 4.1 and 4.4.
@@
Let $\Lambda\subset T^{*}_{\infty}\mathbb{R}^{2}$ be a Legendrian link and choose $\widehat{\Lambda}\subset T^{*}\mathbb{R}^{2}$ to be the union of the Lagrangian cone over $\Lambda$ and all the bounded regions of $\mathbb{R}^{2}\setminus\pi(\Lambda)$, where $\pi:T^{*}M\longrightarrow M$ is the natural projection. For any $U\subset\mathbb{R}^{2}$, we have $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}^{\text{pre}}(T^{*}U)=\operatorname{Sh}_{\widehat{\Lambda}}(U)$. Now, $\operatorname{Sh}_{\widehat{\Lambda}}$ forms a sheaf of categories and the universal property of sheafification implies that $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}(T^{*}U)=\operatorname{Sh}_{\widehat{\Lambda}}(U)$. ((See also [JT24, Appendix B].)) In particular, since $\operatorname{\operatorname{\mu Sh}}_{\widehat{\Lambda}}$ is supported on $\widehat{\Lambda}$, global sections of the Kashiwara–Schapira stack are
@@
where the subscript 0 denotes the full subcategory of sheaves with bounded support in $\mathbb{R}^{2}$. $\Box$